Calculate the change in the standard molar entropy of sulfur trioxide gas, SO3, when it is cooled from a temperature of 100°C to 10°C at a constant pressure of 1 bar.

*use equation ∆S = nCpln(T2/T1)
Answer #1

For SO3 Non lineas Cu= 3R + 3R + (3N-6)R -IWR+).58+ (8X14)-6)R = 3R + GR Cv=qR Cp -c=R CP=Cr+R (p=10R) for n=1 foxs - ncp en/